# BlockFUL: Enabling Unlearning in Blockchained Federated Learning

### v2

@@ Abstract
Index Terms:
machine unlearning, federated learning, dag, blockchain.
@@ I Introduction
I-A Background and Motivations
Various Federated Learning (FL) structures highlight the inherent property of models, where inheritance relationships between models are distinctly visible, including traditional FL [1, 2, 3], multi-layer FL [4, 5], multi-server FL [6, 7], and decentralized FL [8, 9, 10]. These diverse structures, where the intricate connections between models naturally form a directed acyclic graph (DAG), demonstrate the extensive relevance of model inheritance. To enhance the integrity and traceability of these models, Blockchained FL has emerged [11, 12, 13], featuring “certified-by-blockchain” capabilities to ensure that the lineage and updates of each model are verifiable and trustworthy.
@@ I-A Background and Motivations
While helping prevent FL models from being tampered with [14, 15, 16] and provides traceability [17, 18, 19], using blockchains stops the models from being rectified (when needed, e.g., some training data contains sensitive information or is later identified as questionable or contaminated). In addition, any user is entitled to request to eliminate the impact of its personal training data on machine learning models [20, 21]. In this sense, machine unlearning [22, 23, 24, 25, 26] and the capability of editing and updating blocks in Blockchained FL systems [27, 28, 29, 30, 31] become relevant.
@@
An advanced one among existing machine unlearning methods is Sharded, Isolated, Sliced, and Aggregated (SISA) [22], which divides data into independent slices and trains them separately. Only the affected model slice needs to be re-trained to improve unlearning efficiency. Unfortunately, the inheritance relationships among models prevent SISA unlearning from being executed in parallel. This is due to the nature of incremental learning, where knowledge is inherited along with the reference between models. This nature renders SISA incapable of isolating unlearning processes since models need to be broadcast to all clients in any round [32, 33, 34, 35, 36]. With SISA ruled out, alternatives such as re-training [37, 38, 39, 40, 20, 41] and gradient ascent [42, 43, 44] come to the fore. Re-training uses an updated dataset to re-train the entire model. Gradient ascent adjusts the model weights to generate larger errors on known data, thereby reducing its reliance on previously learned content and achieving “forgetfulness”.
@@
Integrating blockchains with FL under model inheritance relations can incur significant overhead in the unlearning processes. One reason is that although primitives, such as the Chameleon Hash (CH), have been considered for conditionally editing blockchains [45, 46] at non-negligible costs with complex operations, historical models are recorded on-chain for certification purposes throughout iterations. Unlearning may require editing multiple related blockchain records to remove just one class of knowledge, which significantly increases the complexity of the unlearning operations. Another reason arises from the model inheritance, where a child model inherits and extends its parent model’s characteristics, structure, and parameters. When performing an unlearning operation, the requested model and all its child models need to be updated. To this end, two critical Research Questions (RQs) need to be addressed.
  • RQ1: How can a Blockchained FL framework unlearn historical models while mitigating the non-negligible costs of frequent blockchain edits and updates to all inherited models in FL systems?
@@
I-B Contributions
This paper presents a novel Blockchained Federated Unlearning (BlockFUL) framework, which ensures model inheritance within generic FL systems and supports effective unlearning without compromising model interrelationships. BlockFUL features a dual-chain architecture—comprising a live chain and an archive chain—to provide users with appropriate access for unlearning. Designed for flexibility, BlockFUL supports a plug-and-play approach. For illustration purposes, primitives such as CH and gradient ascent are considered in this paper. Nevertheless, CH and gradient ascent can be readily substituted with alternative redactable blockchain technologies and unlearning methods, such as re-training.
The contributions of this paper are summarized as follows:
@@ I-B Contributions
    We introduce BlockFUL, the first comprehensive framework enabling unlearning within Blockchained FL. This framework features a dual-chain structure: an archive chain for storing all historical models, and a live chain that securely shares the latest updates and supports flexibility across any redactable blockchain method, including widely used CH-based technologies. BlockFUL mitigates the impact of model updates on inherited models, ensuring the traceability and integrity of FL models during unlearning tasks.
@@
    We introduce two unlearning paradigms specifically considered in BlockFUL: parallel unlearning and sequential unlearning. These approaches facilitate the optimization of multiple model unlearning tasks within Blockchained FL, enabling the simultaneous update of multiple transactions and blocks via a single consensus operation. This strategy reduces the frequency of interactions between the FL and blockchain, and alleviates the burden on consensus operations.
@@
    We comprehensively evaluate the effectiveness of BlockFUL by implementing two prevalent unlearning methods paired with the paradigms: parallel unlearning with gradient ascent, and sequential unlearning with re-training. Our analysis confirms that parallel unlearning effectively reduces dependency on data slated for unlearning and converges reliably. Additionally, we assess the overhead of CH and block update operations and compare the computational and communication costs associated with each method, providing a holistic view of their efficiency and practicality.
@@
Experimental results confirm that sequential unlearning with re-training is highly effective. The model, after unlearning, achieves nearly zero accuracy for data meant to be unlearned and an accuracy of 94.71% for retained data across various models and datasets while maintaining model inheritance. However, this method can be time-intensive for models with deep inheritance chains and requires full participation from all involved clients. In contrast, parallel unlearning with gradient ascent shows varied effectiveness based on different models and inheritance depths—the levels of hierarchical dependencies within the inheritance DAG among models, with results including 1.67% unlearning effectiveness in AlexNet and 77.67% for retained data.
I-C Paper Organization
@@
The rest of this paper is organized as follows. The related works are reviewed in Section II. The proposed BlockFUL framework is provided in Section III. The designed dual-chain architecture is presented in Section IV, followed by two unlearning paradigms in BlockFUL introduced in Section V. Section VI presents the comparative experiments between the gradient ascent and re-training methods implemented in the new BlockFUL framework. Section VII concludes this work.
@@
II Related Work
Existing federated unlearning (FUL) research has been focused primarily on parameter adjustment and re-training processes. For instance, Halimi et al. [20] reversed a learning process by training the model to maximize local empirical losses, and executed deletion of client contributions using Projected Gradient Descent (PGD) at clients. Wu et al. [43] utilized class disassociation learning, client disassociation learning, and sample disassociation learning, and utilized reverse Stochastic Gradient Ascent (SGA) and Elastic Weight Consolidation (EWC) for joint unlearning of these three types of requests. FedEraser [37] utilized the historical parameter updates re-trained by the central server during FL to reconstruct the unlearning model. Wu et al. [47] eliminated client contributions by subtracting accumulated historical updates from the model, and utilized knowledge distillation methods to restore model performance without using client data. FRU [21] eliminates user contributions by rolling back and calibrating historical parameter updates, and then utilizes these updates to accelerate federated recommendation reconstruction. KNOT [48] introduces cluster aggregation and formulates the client clustering problem as a dictionary minimization problem for re-training processes. Liu et al. [41] utilized first-order Taylor expansion approximation techniques to customize a diagonal empirical Fisher information matrix-based fast re-training algorithm. FFMU [49] utilizes nonlinear function analysis techniques to refine local machine unlearning models into output functions of Nemytskii operators, maintaining unlearning quality while improving efficiency. Wang et al. [50] utilized CNN channel pruning to remove information about specific categories from the model for FUL processes.
Existing FUL studies predominantly assess whether the latest model version has been successfully unlearned, often overlooking crucial security challenges in FUL systems, such as trustworthiness, version control, and the traceability and accountability of unlearning iterations. They have not adequately addressed scenarios involving blockchain integration, where the need for tamper-proof data integrity and traceability introduces significant system overhead. These oversights become particularly problematic when integrating certified-by-blockchain with model inheritance across various prevalent FL structures, as this combination significantly amplifies the complexity and cost of FUL processes. To this end, our proposed framework, BlockFUL, is the first to enable unlearning capability in Blockchained FL settings, addressing both the unique challenges of model inheritance and the additional overhead incurred during blockchain operations.
III System Model
In this section, we describe the proposed BlockFUL framework, with unlearning capabilities in Blockchained FL. BlockFUL features a new dual-chain structure—comprising a live chain and an archive chain—to provide users with appropriate access for unlearning. BlockFUL ensures model inheritance within various FL structures and supports effective and efficient unlearning with comparable operations costs. Table I summarizes the notation.
TABLE I: Notation and definition
Notation Definition
$Tx_{i,j}$ The $j$-th transaction of the $i$-th block
$CH\left(i,j\right)$ The CH value in $Tx_{i,j}$
$v(i,j)$ The random number in $Tx_{i,j}$
$URI_{w}\left(i,j\right)$ The model identifier in $Tx_{i,j}$
$R_{Tx}\left(i,j\right)$ A list of transactions referenced by $Tx_{i,j}$
$\mathcal{T}(i,j)$ The payload contains $URI_{w}\left(i,j\right)$ and $R_{Tx}\left(i,j\right)$
$w_{s}$ The starting model of to-be-unlearned data in FL
$\nabla\theta_{s}$ Gradient of $w_{s}$ update
$w_{y}$ An inheritance model of $w_{s}$
$P$ The set of paths from $w_{s}$ to $w_{y}$
$w_{j}$ on $p_{i}$ Traverse the $j$-th model on the $i$-th path except $w_{s}$
$N^{R_{j}}$ The number of models referenced by $w_{j}$
$\nabla\theta^{k}_{s}$ Gradient of the $k$-th starting model $w_{s}$ update
$p^{\nabla\theta^{k}_{s}}_{i}$ The $i$-th path with gradient $\nabla\theta^{k}_{s}$
$P^{\nabla\theta^{k}_{s}}$ The set of paths with gradient $\nabla\theta^{k}_{s}$
$d_{i}$ The depth of path $p_{i}$
$W$ The set of starting updated models
$N(W)$ The set of models that inherit from $W$
$N^{i}(W)$ The set of $i$-hop inherited models of $W$
$C_{tran}$ Transmission cost for uploading or downloading models
$C_{con}$ Consensus cost
$C_{CH}$ CH update cost
III-A Preparation
@@
Registration. Before the users can participate in the BlockFUL task, they need to register in the network. The users are both model publishers and model users.
Key initialization. Each user generates two key pairs, i.e., the conversation key pair ($pk,sk$) and the CH key pair ($CH_{pk},CH_{sk}$).
@@
Key usage and permissions. $pk$ and $sk$ are used for signing and verifying a user’s legitimate identity. The CH private key $CH_{sk}$ is assigned to a committee. As a result, the committee can participate in the redactable process of the blockchain. Meanwhile, the CH public key $CH_{pk}$ is broadcast in the network. The users only have permission to share and use the model, and cannot perform change operations.
III-B Training Models
@@
As shown in Fig. 1, the BlockFUL framework is constructed based on an inheritance structure [51], where multiple users participate in FL training for a task, and each user can train multiple models. Consider a weighted directional graph $G=(V,E)$, where $V=(GV,MV)$. Herein, $GV$ denotes the creation model node from which a task is issued, and $MV=\left\{mv_{1},mv_{2},\ldots,mv_{j}\right\}$ collects model nodes in the network that participate in an FL task. The weighted edges, collected by $E=\left\{e_{01},e_{10},\ldots,e_{xx}\right\}$, represent the links of inheritance relationship between the user models in the network.
@@ III-B3 Training
where $\phi$ denotes the training settings, including the learning rate and batch size; and $\daleth\left(\cdot\right)$ denotes the user’s training function for the task.
III-B4 Evaluation
@@ III-B5 Node generation
III-C Unlearning Models
An unlearning process refers to scenarios where users participating in an FL task, if the training data later triggers privacy concerns or is publicly identified as damaged, a part of the training data needs to be withdrawn according to regulations such as GDPR [52]. Unlearning tasks are employed to remove this data’s influence from the models, ensuring that previously trained models can no longer recognize such data. In BlockFUL, this unlearning process needs to be conducted across the inheritance DAG, followed by an additional blockchain unlearning process.
III-C1 Inheritance unlearning phase
@@
The unlearning request is initiated by a specific node. Employing various unlearning methods, all inherited models influenced by this node need to be updated, as shown by the green lattice nodes in Fig. 1. The updating process continues until the model nodes reach the latest generated nodes. Two cases are studied in Section V.
III-C2 Blockchain unlearning phase
As described in Section 3, we design the block structure in the live chain. Each transaction has a transaction reference list $R_{Tx}$ to record the reference relations of its transactions. Based on these relations, we can obtain the entire inheritance model network structure, which allows all inherited transactions to be updated comprehensively.
III-D Player Model
User. Two assumptions of user behaviors are considered. First, model inheritances are time-sensitive. New users joining the FL task cannot train with models from before a certain block height. Second, all users participating in the FL task are considered honest-but-curious. Users honestly follow the procedures, which grant them access only to the live chain to retrieve the latest model versions. Users are then curious about inferring sensitive training data from these models, which is effectively mitigated by executing the unlearning process.
Adversary. Attackers refer to external malicious entities who do not contribute to the Blockchained FL [53, 54] and could alternatively attempt to disrupt the blockchain consensus by gaining control of more than the faulty threshold from the committee. This risk can be mitigated by prevalent committee election mechanisms powered by reliable pseudo-random generators used in advanced consensus algorithms or sharding technologies [55].
Committee. The committee consists of users elected to validate transactions and maintain the network’s integrity. While members are expected to adhere to protocol specifications to prevent fraud, the system can withstand a fraction of its members being compromised; exceeding this fraction may result in failed validations and security breaches, threatening the reliability and effectiveness of the Blockchained FL process.
III-E Design Goal

### v3

@@ I Introduction
Blockchained Federated Learning (FL) integrates blockchain with FL to enhance model trustworthiness and regulatory compliance, and support tokenized model ownerships with decentralized model governance. Its certified-by-blockchain mechanism ensures that model lineage and updates are verifiable and tamper-proof, enabling secure model sharing and facilitating commercialization. In healthcare, Blockchained FL facilitates disease diagnosis models, extending to applications like personalized treatments and drug development while ensuring traceability and auditability [1]. In finance, it supports financial prediction models and derivative products, maintaining data privacy, transparency, and regulatory adherence [2]. Additionally, Blockchained FL aligns with various FL structures—including traditional [3], multi-layer [4], multi-server [5], and decentralized FL [6]—where model inheritance relationships naturally form a directed acyclic graph (DAG), reinforcing its relevance in large-scale collaborative learning.
@@
While helping prevent FL models from being tampered with [2] and provides traceability [1], using blockchains stops the models from being rectified (when needed, e.g., some training data contains sensitive information or is later identified as questionable or contaminated). In addition, any user is entitled to request to eliminate the impact of its personal training data on machine learning models [7, 8]. In this sense, machine unlearning [9] and the capability of editing and updating blocks in Blockchained FL [10] become relevant.
@@
An advanced one among existing machine unlearning methods is Sharded, Isolated, Sliced, and Aggregated (SISA) [11], which divides data into independent slices and trains them separately. Only the affected model slice needs to be re-trained to improve unlearning efficiency. Unfortunately, the inheritance relationships among models prevent SISA unlearning from being executed in parallel. This is due to the nature of incremental learning, where knowledge is inherited along with the reference between models. This nature renders SISA incapable of isolating unlearning processes since models need to be broadcast to all clients in any round [12]. With SISA ruled out, alternatives such as re-training [13, 7, 14] and gradient ascent [15] come to the fore. Re-training uses an updated dataset to re-train the entire model. Gradient ascent adjusts the model weights to generate larger errors on known data, thereby reducing its reliance on previously learned content and achieving “forgetfulness”.
@@
Integrating blockchains with FL under model inheritance relations can incur significant overhead in the unlearning processes. One reason is that although primitives, such as the Chameleon Hash (CH), have been considered for conditionally editing blockchains [16] at non-negligible costs with complex operations, historical models are recorded on-chain for certification purposes throughout iterations. Unlearning may require editing multiple related blockchain records to remove just one class of knowledge, which significantly increases the complexity of the unlearning operations. Another reason arises from the model inheritance, where a child model inherits and extends its parent model’s characteristics, structure, and parameters. When performing an unlearning operation, the requested model and all its child models need to be updated.
To this end, two critical Research Questions (RQs) need to be addressed.
  • $\bullet$
    RQ1: How can a Blockchained FL framework unlearn historical models along with all inherited models while maintaining the traceability and integrity of models?
  • $\bullet$
@@
Contributions. This paper presents a novel Blockchained Federated Unlearning (BlockFUL) framework, which ensures model inheritance within generic FL systems and supports effective unlearning without compromising model interrelationships. BlockFUL features a dual-chain architecture—comprising a live chain and an archive chain—to provide users with appropriate access for unlearning. Designed for flexibility, BlockFUL supports a plug-and-play approach.
The key contributions of this paper are as follows:
@@
    We propose the BlockFUL framework, the first comprehensive framework enabling unlearning within Blockchained FL. BlockFUL adopts a dual-chain architecture: a live chain and an archive chain. This architecture ensures the traceability, compliance, and integrity of model updates.
@@
    To embody and operationalize BlockFUL, we design two distinct unlearning paradigms: Parallel and sequential unlearning. Parallel unlearning enables concurrent updating of multiple models using a single consensus, significantly reducing consensus frequency and computational overhead. Sequential unlearning serializes unlearning operations to prioritize the precision and order of model revisions.
@@
    We validate the feasibility and practicality of BlockFUL for both paradigms by demonstrating its compatibility with existing primitives, including CH-based redactable blockchain [16], gradient ascent unlearning [15] and re-training unlearning [13]. We also analytically prove the convergence of the unlearning processes and quantify the computation, communication, and consensus costs for both parallel and sequential paradigms.
@@
Experimental results confirm that sequential unlearning with re-training is highly effective, achieving nearly zero accuracy for unlearned data and 94.71% accuracy for retained data across various models and datasets while preserving model inheritance. However, it can be time-intensive for deep inheritance chains and requires full client participation. In contrast, parallel unlearning with gradient ascent shows varying effectiveness depending on model type and inheritance depth, with unlearning effectiveness ranging from 1.67% in AlexNet to 77.67% for retained data.
@@
The rest of this paper is organized as follows. The related works are reviewed in Section II. The preliminary in Section III. The proposed BlockFUL framework is provided in Section IV. The designed dual-chain architecture is presented in Section V, followed by two unlearning paradigms in BlockFUL introduced in Section VI. Section VII presents the comparative experiments between the gradient ascent and re-training methods implemented in the new BlockFUL framework. Section VIII concludes this work.
II Related Work
II-A Federated Unlearning
Existing federated unlearning (FUL) studies focus primarily on parameter adjustment and re-training processes.
II-A1 Parameter adjustment
Halimi et al. [7] reversed a learning process by training the model to maximize local empirical losses and executed deletion of client contributions using Projected Gradient Descent (PGD) at clients. Wu et al. [17] utilized class disassociation learning, client disassociation learning, and sample disassociation learning, employing reverse Stochastic Gradient Ascent (SGA) and Elastic Weight Consolidation (EWC) for joint unlearning of these three types of requests. Wu et al. [18] eliminated client contributions by subtracting accumulated historical updates from the model and utilized knowledge distillation methods to restore model performance without using client data. FRU [8] eliminates user contributions by rolling back and calibrating historical parameter updates, and then utilizes these updates to accelerate federated recommendation reconstruction. FFMU [19] utilizes nonlinear function analysis techniques to refine local machine unlearning models into output functions of Nemytskii operators, maintaining unlearning quality while improving efficiency. Wang et al. [20] utilized CNN channel pruning to remove information about specific categories from the model for FUL processes.
II-A2 Re-training
FedEraser [13] utilized the historical parameter updates re-trained by the central server during FL to reconstruct the unlearning model. KNOT [21] introduces cluster aggregation and formulates the client clustering problem as a dictionary minimization problem for re-training processes. Liu et al. [14] utilized first-order Taylor expansion approximation techniques to customize a diagonal empirical Fisher information matrix-based fast re-training algorithm.
Existing FUL studies predominantly assess whether the latest model version has been successfully unlearned, often overlooking crucial security challenges in FUL systems, such as trustworthiness, version control, and the traceability and accountability of unlearning iterations. They have not adequately addressed scenarios involving blockchain integration, where the need for tamper-proof data integrity and traceability introduces significant system overhead. These oversights become particularly problematic when integrating certified-by-blockchain with model inheritance across various prevalent FL structures, as this combination significantly amplifies the complexity and cost of FUL processes.
II-B Blockchain Federated Unlearning
Islam et al. [22] utilized redactable blockchains for flexible data updates in smart consumer electronics, enabling model update recording while allowing historical data modifications. However, they did not address the traceability and auditability of unlearning models. To enhance trustworthiness, Lin et al. [23] proposed a federated unlearning proof protocol using the CH function, allowing target clients to invoke an unlearning rewrite function to erase model updates and associated data. Zuo et al. [24] introduced differential privacy to ensure transparency while preserving data confidentiality and reducing blockchain computational overhead. Rather than modifying on-chain data, their approach records a complete historical trace, leveraging blockchain immutability.
Different from these existing studies, our blockchain system features a dual-chain structure that supports redactability while ensuring model traceability and auditability. It also addresses real-world application requirements by enabling parallel model updates and the simultaneous update of multiple starting models. By judiciously optimizing blockchain operations, we minimize the overhead associated with redactability.
III Preliminary
III-A Federated Learning
The goal of FL is to minimize the weighted average loss across $K$ clients, each with its own local dataset. The Federated Averaging (FedAvg) algorithm is typically used for this process [3].
  • $\bullet$
    Local Training: Each client performs several gradient descent updates on its local dataset to optimize its model.
  • $\bullet$
    Model Aggregation: The server collects the local model parameters from all $K$ clients and computes their weighted average to update the global model.
  • $\bullet$
    Model Distribution: The server sends the updated global model parameters to all clients for the next training round.
III-B Chameleon Hash
The CH function allows modifying message data and its random value without changing the hash, provided the private key is known. This feature is especially useful in blockchain for scenarios requiring data adjustments [10].
  • $\bullet$
    CH Generation: Generate a pair of public and private keys. The public key $pk$ is used for hash computation, while the private key $sk$ (trapdoor) is used for subsequent hash value modification. Given a message $m$ and a random number $r$, the CH value $h$ is computed using the CH function $CH$, resulting in $h=CH\left(pk,m,r\right)$.
  • $\bullet$
    CH Verification: For a given message $m$, random number $r$ and hash value $h$ the verification process is performed by checking whether $h=CH\left(pk,m,r\right)$. If this equation holds, the verification function $Verify(pk,m,r,h)$ returns $True$, confirming the validity of the hash value.
  • $\bullet$
    CH Update: When in possession of the private key $sk$, the hash value can be updated. Given the original message $m$, the original random number $r$, a new message $m^{\prime}$, and a new random number $r^{\prime}$, the updating process is carried out using the updating function $Update$, which computes $r^{\prime}=Update\left(sk,m,r,m^{\prime}\right)$. This ensures that the new hash value $h^{\prime}=CH\left(pk,m^{\prime},r^{\prime}\right)$ remains identical to the original hash value $h$, satisfying $h=h^{\prime}.$
III-C Machine Unlearning
An unlearning process refers to scenarios where users participate in an FL task, if the training data later triggers privacy concerns or is publicly identified as damaged, a part of the training data needs to be withdrawn according to regulations, such as GDPR [25]. Unlearning tasks are used to remove the influence of these data from the models, ensuring that previously trained models cannot recognize such data [11]. The unlearning process is typically achieved through two primary methods: Re-training and gradient ascent.
  • $\bullet$
    Re-training: This straightforward approach to unlearning involves removing the specific data from the training set, re-training the model with the remaining data, and ensuring the new model excludes the unlearned information [13].
  • $\bullet$
    Gradient ascent: This approach reverses optimization to increase the error on specific data, eliminating its influence by computing the loss, applying gradient ascent to maximize it, and iteratively reducing the dependence of the data [15].
IV System Model
This section presents the proposed BlockFUL framework, enabling unlearning in Blockchained FL. BlockFUL introduces a dual-chain structure—comprising a live chain and an archive chain—to facilitate controlled unlearning access. It preserves model inheritance across FL structures while ensuring efficient unlearning with comparable operational costs.
@@ IV System Model
IV-A Preparation
IV-A1 Registration
@@
Before the users can participate in the BlockFUL task, they need to register in the network. The users are both model publishers and model users.
IV-A2 Key initialization
Each user generates two key pairs, i.e., the conversation key pair ($pk,sk$) and the CH key pair ($CH_{pk},CH_{sk}$).
IV-A3 Key usage and permissions
@@
pkpk and $sk$ are used for signing and verifying a user’s legitimate identity. The CH private key $CH_{sk}$ is assigned to a committee. As a result, the committee can participate in the redactable process of the blockchain. Meanwhile, the CH public key $CH_{pk}$ is broadcast in the network. The users only have permission to share and use the model, and cannot perform change operations.
IV-B Training Models
@@
As shown in Fig. 1, the BlockFUL framework is constructed based on an inheritance structure [6], where multiple users participate in FL, and each user can train multiple models. Consider a weighted directional graph $G=(V,E)$, where $V=(GV,MV)$. Herein, $GV$ denotes the creation model node from which a task is issued, and $MV=\left\{mv_{1},mv_{2},\ldots,mv_{j}\right\}$ collects model nodes in the network that participate in an FL task. The weighted edges, collected by $E=\left\{e_{01},e_{10},\ldots,e_{xx}\right\}$, represent the links of inheritance relationship between the user models in the network.
@@ IV-B3 Training
[…]
IV-B4 Evaluation
@@ IV-B5 Node generation
IV-C Unlearning Models
In BlockFUL, this unlearning process needs to be conducted across the inheritance DAG, followed by an additional blockchain unlearning process.
IV-C1 Inheritance unlearning phase
@@
The unlearning request is initiated by a specific node. Employing various unlearning methods, all inherited models influenced by this node need to be updated, as shown by the green lattice nodes in Fig. 1. The updating process continues until the model nodes reach the latest generated nodes. Two cases are studied in Section VI.
IV-C2 Blockchain unlearning phase
As described in Section V-A, we incorporate CH into the live chain’s block structure to enable editability. Each transaction contains a CH value and a reference list $R_{Tx}$, forming a complete inherited model network and allowing updates to inherited transactions without disrupting the blockchain structure.
IV-D Trust Model and Security Analysis
IV-D1 System entities
We consider three types of participants in the system.
  • $\bullet$
    User. Users in the FL task are assumed to be honest-but-curious, complying with protocols and accessing models only on the live chain. Model inheritance is time-sensitive; new users cannot access models published before a specific block height, which prevents retrospective inference of early training data at the system level. BlockFUL can also employ $(\epsilon,\delta)$-differential privacy [26] to prevent dishonest users or attackers from inferring other users’ training data through the models.
  • $\bullet$
    Committee. Committee members are elected blockchain miners responsible for achieving consensus on both the live chain and archive chain. A two-stage committee selection process, including pool formation and member selection, can select responsible committee members.
    1. 1.
      In the Initial Pool Formation stage, BlockFUL can leverage a Proof-of-Stake (PoS) mechanism [27], which requires candidates to commit substantial stakes to join the committee pool. The pool formation can be augmented with a dynamic trust scoring scheme [28] which continuously evaluates member behavior and excludes those who fail to maintain required trust thresholds.
    2. 2.
      In the Member Selection stage, committee members are chosen from the qualified pool through a Verifiable Random Function (VRF)-based selection protocol [29], which ensures both unbiased randomness and cryptographic verifiability while maintaining complete transparency in the selection process. The two-stage selection process enhances system redundancy and attack resistance. The PoS and trust score schemes work together to regulate committee behavior, ensuring committee members conduct responsible blockchained FL and FUL consensus while maintaining an up-to-date view of the live chain. Any misconduct by malicious committee members results in their exclusion from the committee pool and forfeiture of their deposited stake.
  • $\bullet$
    Adversary. Adversaries are malicious entities that aim to disrupt blockchain consensus processes, including model inheritance and update verification procedures.
IV-D2 Quantitative security analysis
Let $P$ denote the size of the committee pool containing $M$ malicious candidates. Using VRF, $N$ committee members are randomly selected from the pool. Given that the trust score mechanism disincentivizes malicious behavior, we model each malicious committee member as attacking with probability $\rho$. BlockFUL maintains consensus integrity as long as the number of malicious votes remains below the consensus protocol’s fault tolerance threshold $f$, where $f<M\leq N$. The attack success rate, denoted by $\mathcal{A}$, can be given by
$\mathcal{A}=\sum_{k=f+1}^{M}\left(\frac{{\binom{M}{k}\binom{P-M}{N-k}}}{{\binom{P}{N}}}\times\sum_{i=f+1}^{k}\binom{k}{i}\rho^{i}(1-\rho)^{k-i}\right)$ (4)
Take a Practical Byzantine Fault Tolerance (PBFT) consensus protocol tolerating up to $f=\left\lfloor\frac{N-1}{3}\right\rfloor$ malicious members for example. Setting $P=30$, $M=10$, $N=21$, and $\rho=0.2$, we obtain $f=6$ and $\mathcal{A}=0.0000625$, demonstrating the system’s robust security against potential compromises.
IV-E Design Goal